OpenStudy (anonymous):

Find out whether the function converges or dierges, and if the function converges, then integrate the given equation integrate from 0 to +infinity sin^2(x) / (x^6+1) dx

OpenStudy (anonymous):

\[\int\limits_{0}^{\infty} \sin^2x/(x^6+1) dx\]

OpenStudy (anonymous):

Is this problem from a complex analysis course?

OpenStudy (anonymous):

That is what Mathematica gave \[ \int_0^{\infty } \frac{\sin ^2(x)}{x^6+1} \, dx=-\frac{\pi \left(1+e \left(-2 e+\sqrt{3} \sin \left(\sqrt{3}\right)+\cos \left(\sqrt{3}\right)\right)\right)}{12 e^2} \]